OpenStudy (iwanttogotostanford):

ACT MATH REVIEW PLEASE HELP: what is the slope of any parallel line to the line 7x+9y=6?

OpenStudy (iwanttogotostanford):

@Vocaloid

Vocaloid (vocaloid):

well, first calculate the slope, do you know how do to that?

OpenStudy (iwanttogotostanford):

y2-y1/x2-x1

Vocaloid (vocaloid):

uh, well, that would work if we were given two points, but we don't have that here, so we're going to use a different method

Vocaloid (vocaloid):

are you familiar with the equation y = mx + b?

OpenStudy (iwanttogotostanford):

yes

Vocaloid (vocaloid):

well, we're going to take our equation 7x + 9y = 6 and re-arrange it so it looks like y = mx + b so let's start off with 7x + 9y = 6 subtract 7x from both sides, what do you get?

OpenStudy (iwanttogotostanford):

9y=-1x

Vocaloid (vocaloid):

uh, not quite remember, you can't combine 7x and 6, so we leave it like this: 9y = 6 - 7x, with me so far?